OpenStudy (anonymous):

Series: Determine the limit of the following: summation from m=2 to infinity of (arctan1)^m

OpenStudy (zarkon):

this is just a geometric sum...where are you stuck? also, I don't believe you wrote the question correctly...what limit are you talking about...do you mean this \[\sum_{m=2}^{\infty}(\arctan(1))^m=\lim_{n\to \infty}\sum_{m=2}^{n}(\arctan(1))^m\] a little hint: \(\displaystyle \arctan(1)=\frac{\pi}{4}\)